Derek wants to withdraw $10,044.00 from his account 7.00 years from today and $13,822.00 from his account 13.00 years from today. He currently has $2,966.00 in the account. How much must he deposit each year for the next 13.0 years? Assume a 6.72% interest rate. His account must equal zero by year 13.0 but may be negative prior to that.

Please show steps (with calculator)
Answer #1

I/Y=6.72%
N=7
PMT=0
FV=10044
CPT PV=6370.68

I/Y=6.72%
N=13
PMT=0
FV=13822
CPT PV=5934.38

I/Y=6.72%
N=13
PV=6370.68+5934.38-2966
FV=0
CPT PMT=1099.76

Hence, deposit each year=1099.76